North Eastern Locomotive Preservation Group

  • Increase font size
  • Default font size
  • Decrease font size
Q6 Blog

A record of events, activities and our work with the Q6.

As the Covid-19 pandemic had reduced the requirement for locomotives on the North Yorkshire Moors Railway, in June 2020 a small group of volunteers, working within the Covid-19 protocol, made an early start on the plan to lift the boiler out of the frames. This was to allow an external examination of the boiler by the insurance company, and to allow the boiler ticket to be extended to 2028 - 10 years from the date of the internal examination. The opportunity was taken at this time to service the axle boxes, horns and related "bottom end" components.

The locomotive has had an "internal" steam test which has revealed a small number of snags to be rectified. Afer the formal steam test, witnessed by our insurance company boiler inspector, the locomotive should be ready for service again.

Q6 10th 12th 13th and 14th April 2021

Another multi day report from Ian Pearson, plus some admin requests from MPD management at the end.

On Saturday 10 April, Bill Dobson, Ian Pearson and Nigel Hall were present to move three locomotives so as to give some space to refurbish the Q6 cab. No 7 Road has a stop board on it and is in very poor condition, so shunting on it has been stopped. Bill had borrowed a Tirfor winch from the MPD instead of having to use pinch bars. First job was to remove various items from the north end of No 8 Road. This included a velocipede which is now in the Armstrong Oilers building. Then the Bullied Pacific Hartland's tender (with Hartland attached) was winched back to create space at the front of No 8 Road to fit the Q6 cab (see photo). The B1 was then moved outside onto No 7 Road with both a struggle and help from others working on the K1 and J27. The Q6 was then moved to the front of No 7 Road to facilitate the cab removal. Nigel Hall continued to undercoat various cladding items during the afternoon.

Bill Dobson winching Hartland's tender to make room for work on 63395's cab in Deviation Shed on 10 April 2021 -Ian Pearson

Bill went in on the Monday to see if a shunt could take place to continue moving the B1 and Q6 to allow the Q6 cab to be removed. This went ahead and the cab was removed with the JCB front loader and set down on No 8 Road.

Tuesday was a fine, frosty, sunny, morning which warmed up later with Bill Dobson, Nigel Hall, Jon Bradley, Trevor Wilford and Ian Pearson present. The main job was to move the cab into Deviation Shed on No 8 Road, then turn it so the cab roof was vertical and easier to clean, prep and varnish. Trevor spent most of the day making up the new stainless steel cotters which will secure the ash pan to the firebox. He then connected the new spool of mig wire to the welding set. Meanwhile, moving the cab took the remaining four nearly all day, using both pull lifts and a system of rollers using old boiler tubes. Nigel did get some cladding painting done later on in the afternoon however.

Wednesday was another sunny, frosty morning, again warming up by late morning. Present were Bill Dobson, Nigel Hall, Ed Bolam and Ian Pearson, along with Terry Newman and myself. Ian, Ed and Nigel had an NYMR coronavirus lateral test: all were negative. Terry came in to do some special soldering of the new copper pipe work for the axle box lubrication system, and also soldered some pipe work to fittings to a pressure gauge to check that the lubricators are delivering oil. Bill was assisting and advising Terry with the copper pipe soldering. With the new pressure gauge fitted to the valve and cylinder lubricator, a test was carried out successfully. A start was made on the axle box lubricator but more tests are required.

I was on my first visit to Deviation Shed for a year, and had various business items to discuss with shed management and others. As I had also had a birthday in the 12 months since my last visit, I had brought in a chocolate cake which was shared at lunch time. I also made afternoon tea and gave Nigel Hall a hand in the morning to move some of the cladding sheets to make space for Nigel and Ed, with a little assistance from Ian, to later lift into position the cladding sheets that Ed had been dressing during the day, ready for Nigel to paint the insides. Nigel had also started on cleaning and prepping the inside cab roof for revarnishing (see photo). Ian took out six bottom doors from the boiler: these are in a box in the workshop, then ran a 1/2 “ die nut down four studs on the cab blower valve.

63395 cab roof being rubbed down by Nigel Hall on 14 April 2021 - Ed Bolam


Q6 - 4/6/7 April 2021

Ian reports that, on Sunday (4 April), Bill and he went to Deviation Shed to move all the boiler cladding from the centre to the front area of of No 6 Road. They also moved various other items that were stored in the four foot of No 6 Road to the side footpath. That then allowed No. 5 locomotive to be moved back to buffer up to 80135 using pinch bars, helped by Richard De Sadeleer who was working with Paul Hutchinson that day. After a break for lunch, a bench was then set up inside No 6 Road, onto which some painted sheets were moved, before setting up the outside benches with fresh cladding sheets for dressing and priming.

For the regular working party on Tuesday, it was fine, cold, and sunny with hail and snow showers during the afternoon - quite a contrast with last week! Attending for the day were Bill Dobson, Nigel Hall, Jon Bradley, Nigel Bill, Trevor Wilford, Gordon Wells and Ian himself. After a quick cuppa, Trevor was again working mainly on the ash pan, where extra welding was required on the rear inside of the ash pan, and also annealing two sections of copper pipe. Nigel Hall and Jon Bradley were equally working all day dressing and priming the cladding sheets. Nigel Bill spent most of the day drilling out larger bracket holes to suit our bolts. These brackets were newly delivered for clamping the copper pipe work. Barry Nesom brought in the new lubrication blocks which he had made down at the MPD machine shop. These are to be attached to the top of each axle box and then connected to the oil supply. Gordon worked all day on the copper pipe work and we now have what looks like the new axle box lubricator system (see photo).

Ian and Bill removed all the grate from the firebox. The front bars have been stacked on the racks at the side of No 4 Road. The rear set are being kept close by as they will be refitted when the boiler is returned to the frames. Later Bill was working on the axle box fittings.

After lunch, Ian spent all afternoon trying to remove 3 remnants of cotter pins from the foundation ring pegs which got overlooked when the boiler was lifted last year. Unfortunately none would move, even after heating and hammering with a chisel. It was decided it would have to be a drill out job with the cotters being made of stainless steel.

Wednesday was another very cold, fine, day but without the hail and snow showers. With Bill Dobson, Ed Bolam, Nigel Hall, and Ian, but also Mike Bloomfield, available from the K1 team. Nigel Hall and Mike spent the day dressing and priming the cladding sheets, while Ed Bolam spent half of the day drilling and tapping small holes between the frames for the oil pipe brackets. Meanwhile, Bill was busy drilling and tapping the 8 axle box top plugs which were only threaded halfway down.

63395 new lubrication system on 7 April 2021 - Ian Pearson

Ian had acquired 2 new cobalt drill bits from Screwfix on Tuesday evening on his way home to drill through the stainless steel cotter remnants on the foundation ring pegs. Paul Hutchinson made up a long bar with a chisel end, although he did comment that Ian should have been using a heavier hammer, but, as Ian pointed out, a bit awkward when you are laying on the ground to get a swipe at it.

After lunch Ed worked with Ian down at the firebox, and, after a lot of drilling holes into the remnants, they managed to knock the offending pieces out ( see photo with offending pieces removed). The reel of MIG welding wire was finished on Wednesday and Trevor is going to obtain another one.

63395 cotters finally removed from foundation ring pegs on 7 April 2021 - Ian Pearson


Q6 - 30th to 31st March 2021

With the recent minor heatwave, Ian has gone all poetic, and reports that on a beautiful Spring Tuesday, with the hawthorn hedges turning Into a wonderful show of fresh greenery, the team at Grosmont that day was Bill Dobson, Nigel Bill, Trevor Wilford, Jon Bradley, Nigel Hall, Gordon Wells and Ian himself. When Jon arrived he brought the new piston ring replacement and the atomisers which have had new fittings fitted, whilst Nigel Bill delivered some consumables, ie paper towels, gloves, hand cleanser and some hacksaw blades.

There was quite a lot going on, with Trevor working on the ash pan sprinkler system all day. This part of the job is almost completed, which included some profile burning and welding (see photo).Jon and Nigel Hall made up some benches outside Deviation Shed on No 7 Road to dress and paint the cladding (see photo). Bill and Ian removed the left hand piston which required the new replacement ring once Bill had cut the ring.Gordon carried on with the copper pipe work, chamfering the ends of the horn guide pipes so as to allow the oil to drip in the correct place. Nigel Bill completed the work on the gauge glass protectors, fitting new rods and springs which hold the backplate tightly to the glass. He then drilled a row of holes in the ash pan sprinkler pipe.

63395 ashpan showing welding repairs on 30 March 2021 - Ian Pearson

63395 boiler cladding being painted by Nigel Hall on 30 March 2021 - Ian Pearson

Lunch was held alfresco on the picnic tables overlooking the coal heap in lovely warm sunshine. After lunch, Bill and Jon fitted the new piston ring to the left hand piston head and then fitted the piston back into the cylinder and connected it to the crosshead.Ian assisted Trevor all afternoon with dressing and welding jobs on the ash pan.

Wednesday was another fine warm day at Grosmont with Bill Dobson, Ed Bolam, Nigel Hall and Ian Pearson. All attended the coronavirus lateral flow testing being carried out by the NYMR, and all tested negative.

Afterwards Ed set about running a 5/16 tap down all the crinoline holes on the boiler and two section of crinoline stored on No 6 Road - about

150 holes in all. Bill and Ian did a spray test on the ash pan which showed it to be working properly. Ian then fitted four new securing pins to the rocking grate on the ash pan, replacing four nuts and bolts (see photo). Nigel Hall spent all day dressing and priming the cladding sheets, and Bill joined him after lunch. lan fitted wooden dowel plugs to the top of the axle boxes where trimmings were fitted to lubricate the horn guides.

63395 ashpan showing new securing bolt temporarily fitted on 31 March 2021  Ian Pearson

Just before lunch there was a delivery of 9 boxes of boiler insulation which are now stored in Deviation Shed. Lunch was had alfresco again due to the fine weather.


21st - 24th March 2021

With the arrival of Spring, with Daffodils and Primroses appearing by the roadside, and being blessed with some fine sunny weather, Ian has reported on four days of Q6 activity at Grosmont this week.

Sunday 21st.

With the Q6 boiler insurance test booked with Bureau Veritas inspector Glynn Coxall for 2pm on Monday 22nd March, Bill and Ian arranged with MPD staff to put a warming fire in the Q6 on the Sunday afternoon. They both arrived about 2pm and set about cleaning the grate, then setting a warming fire. Bill also carried out some work cutting a slot on the running plate to allow the drive from the new lubricator to be connected to the drive from the cylinder lubricator. Ian fitted the pressure gauge and gauge glass protectors, then lit the warming fire.

Monday 22nd.

Another fine sunny day but on a very cold morning, Ian went in around 6am to get the fire box ash cleaned and lit up again. While waiting for the steam to rise he cleaned the bottom smokebox door hinge strap as it was very rusty. Bill and Gordon came in later in the morning to continue working on the lubricator and fitting new pipe work from the brass oil boxes which will supply oil to the horn guides (see photo). Although the boiler was lit early, it seemed reluctant to make steam, so a polythene sheet was wrapped around the bottom side of the smokebox with sticky tape over the smaller holes around the smokebox. This helped tremendously to draw the fire with the blower in action. By 2pm there was 100 lbs of steam on the gauge, but, fortunately, it was still rising as Glynn had just arrived. So Bill made him a cuppa at Deviation Shed and sorted out the paper work. Then, by 2.20 pm it was not far from blowing off when Glynn came down from Deviation Shed with Bill and carried out his inspection. The safety valves are blowing off slightly light, but everything was satisfactory so the out of the frames test was satisfactorily completed at long last, with a ticket to 2028 subject to annual inspections (see photo).

63395 on out of frames insurance steam test at Grosmont on 22 March 2021 - Ian Pearson

Tuesday 23rd

A fine, slightly overcast, day with bright periods for Bill Dobson, Ian Pearson, Nigel Bill, Gordon Wells and Trevor Wilford. Bill continued with the mechanised fittings on the new lubricator while Gordon continued fitting and completed the rest of the copper pipe work from the brass oil boxes to the horn guides. Nigel cleaned the rust from the top smokebox door hinge strap and greased up both of them, and, after Ian had removed the pressure gauge and gauge glass protectors, Nigel turned his attention to cleaning these. The pressure gauge was then returned to the MPD stores for safe keeping. Trevor put a bend into the steel driving shaft for the new lubricator on Bill's instructions. He then spent the rest of the day working on the ash pan. Ian cleaned all the old ash from the firebox and then started moving it to the ash compound before cleared some water and air hoses which had been used during the steam test.

63395 new lubrication pipes for horn guides on 23 March 2021 - Ian Pearson

Wednesday 24th

Another bright day with Ed Bolam, Bill Dobson, Ian Pearson, Nigel Bill and Terry Newman. Terry came in to solder the new copper pipes that Gordon fitted over the last two days: these are now fully connected to the brass boxes. Nigel Hall called in briefly to see what paint was required for the Q6 cladding repaint.

Ed assisted Ian in removing the balance of ash from underneath the fire grate. Then Ed had a crash course on the big radial drill with Chris Kelley setting it up. Ed then drilled out a 1”hole in seven 4 inch plates which will fit around the pegs on the bottom of the ash pan and then be pinned with a cotter when the ash pan is attached back onto the boiler. After noticing that the rocking grate linkage was bolted with double nuts instead of being pinned (which uses a washer and split pin), and discussion with Nick Simpson, Nigel was asked to give Nick a drawing of what was required and the MPD will have four new pins made. This will rectify a mistake when the MPD altered the grate from fixed fire bars to a rocking grate. Nigel also freed both damper doors with a little gentle persuasion and some oil. Later he made two long pins for the gauge glass protectors, and Chris Cubitt tapped a thread on the bottom end of each one. Nigel is going to acquire some small nuts to fit these and he will then be able to fit new springs. Then the back plate will stay closed. Nigel also removed, cleaned and ground in two steam oil supply test cocks and refitted them back to the supply pipes.

Ian set about removing a piece of backplate which was needed to support the pressure gauge during the steam test. A simple sounding task, but it took all afternoon with some help from Ed. First the gauge glasses had to be removed, then the blower delivery pipe which has air ministry joints, and eventually the section of backplate could be removed. On inspecting the blower valve bottom studs these looked to be wasting away with not much thread left, so they will have to be removed and replaced. The next job was a very dirty one in the smokebox, removing the blower pipe. This was taken with the cab blower pipe back to Deviation workshop for annealing. Bill borrowed the mag drill from the MPD, set it up on the running plate, and drilled the four holes required and then secured the new lubricator in place.

Bill has ordered the boiler insulation but not as much as first calculated due to the fact that about 10 rolls were salvaged from the MPD rubbish skip. Some had got slightly damp on the corners but on inspection they were found to still be suitable for use. The Geoff Drury training still pays dividends.


16th and 17th March 2021

Ian Pearson reports some early morning rain but becoming brighter later before lunchtime on Tuesday. Bill Dobson, Nigel Bill, Jon Bradley, Gordon Wells, Trevor Wilford and Ian Pearson were present, and found, on arrival at the MPD, the Q6 boiler in full steam with Adrian Dennis in charge and about to set the safety valves. A report later in the day from Adrian showed eight minor steam leaks needing attention, and that the safety valves had been satisfactorily set .

Up at Deviation Shed, Nigel lit the stove and then went down to the MPD for some water. Our supply to Deviation Shed had been turned off due to repairs being necessary to the main supply pipe. By late afternoon however, after some new pipe work was fitted, the supply was reinstated.

Jon arrived from Morpeth with the new lefthand piston rings, repaired lubricator and some items for the J72 for Neal Woods attention. After unloading these items, and a welcome cup of tea, Jon gave the inside of the cab another coat of undercoat. The wooden interior from the cab was loaded up into his car up to take home for cleaning and varnishing.

Bill, on receiving the piston rings, found the rear ring fitted OK, but the front one was slightly too tight and he spent a lot of time dressing it to fit. Nigel spent some time on the band saw cutting some steel, making seven 4”x3” plates which will be required when fixing the ash pan back to the boiler.

63395 new lubrication pipework to left hand rear hornguides on 16 March 2021 - Ian Pearson

Gordon spent all day making new copper pipe brackets and fitting them to supply oil to the rear axle box horn guides (see photo of lefthand rear). Nigel also assisted Gordon in removing two old oil pipes that used to feed oil to the right hand side rear axle box . New pipes were then fitted but are not quite finished.

About 11am, Neil Smedley called in on his way to Loftus for his PTS medical, following prior agreement with Ian, to deliver a 15” adjustable spanner, retrieved from a scrap metal bin at a charity shop, and for which many thanks are due to Neil and his eagle eyed son.

Trevor arrived after lunch and continued working on the ash pan sprinkler system, welding brackets and fitting the sprinkler pipe. After lunch it was decided to fit the piston rings and put the piston into the cylinder using the engine hoist and lifting table (see photos).

63395 left hand cylinder and piston rings awaiting fitting on 16 March 2021 - Ian Pearson

Eventually the piston was fitted back into the cylinder, then it was time for home.


63395 left hand piston on hydraulic lifting platform about to be installed on 16 March 2021 - Ian Pearson

Wednesday 17th March 2021 started off with a cloudy, slightly damp, morning, but the sun came out before lunchtime for the much smaller team of Bill Dobson, Ed Bolam and Ian Pearson.

After Ed had got the stove fired up, he spent some time down at the MPD cutting a piece of steel 1 metre in length by 40 x 25mm into 8 pieces, 125mm long, to make a fitting to deliver oil to each axle box from the lubricator. These are to be machined by Barry Neesom next week.

Overnight, Bill had a suspicion that the lefthand front piston ring was still a bit tight. On investigation, it was found that the front ring had dropped into the port. It took some time to get the ring free, but, with the use of a shim, the piston was brought forward and outside the cylinder. The ring was to be removed and dressed some more, but, unfortunately, while trying to remove it, the ring was broken. A new one has been ordered from Matthew Storey Engineering.

After lunch an air line was set up and the vacuum pipe was blown through to remove any scale and dust that may have been inside the pipe after straightening. Also, all the copper pipes have been blown through.

Ed Bolam took the J72 bolts to deliver to Hopetown, and Neal Woods has been advised.

Paul and Angie were busy working on the K1 on Tuesday and Wednesday with assistance from Chris Henwood and Mike Bloomfield.


9th, 10th and 11th March 2021

A three day report from Ian this week.

Tuesday was mild and overcast with some bright periods later, with Bill Dobson, Nigel Bill, Jon Bradley, Trevor Wilford, Gordon Wells and Ian Pearson forming the crew. Ian lit the stove and Nigel got the water supply turned on after it had been off over the weekend. Kettle filled, tea made and then disaster. Checked the biscuit tin to find it was empty! The weekend crew must have eaten them all!

The boiler had the latest two replacement tea tray studs fitted, the bottom right door refitted, and the boiler filled on Monday. Bill then fitted the ejector supply valve on Tuesday morning. Ian cleaned the boiler grate of old ash and prepared a warming fire which was lit early afternoon with the intention of a steam test the next day. Gordon and Bill then worked on the copper pipe, cutting and bending to fit the rear axle boxes, while Nigel painted the brackets with quick drying zinc undercoat and a coat of red gloss later on in the day. Jon painted the five sections of cab backplate with black undercoat and also cleaned the lower cab area with white spirit, before giving the area a coat of black primer. Trevor spent the day working on the ash pan sprinkler system under Bill's guidance. During the afternoon Bill drilled and tapped some of the holes for the oil pipe brackets.

In contrast, Wednesday was a grey mild day, blustery and with heavy rain after lunch, to welcome Bill Dobson, Ed Bolam, Nigel Bill, Terry Newman, Ian Pearson and Jon Bradley (for two hours only). Ed lit the stove while Jon made a cuppa, but, most importantly, the biscuit tin was replenished with fresh packets of biscuits. The Q6 boiler had been relit early in the morning by the MPD boilersmith, and when the NELPG team arrived, the fire was burning nicely but no steam was registered on the gauge except for a wisp from the whistle.

Terry spent all day soldering brass fittings onto the copper pipes. Ed made cotter pins for the left and right trailing gradient pin taper pins and then fitted them. Ian, while looking after the boiler in steam, took off the righthand little end oil pot and fitted a new plug tail trimming after soaking in oil for 24 hours. This was tested for oil flow and later refitted by Ed. Nigel gave the oil pipe brackets another coat of red gloss paint and also completed the monthly check on the Defibrillator. Bill finished drilling the bracket holes and then continued to tap the threads ready for fitting the brackets. There is still another hole to drill, but this won’t be done until the pipe work is in the correct position as the bracket may need to be swivelled slightly. Jon painted the five sections of cab backplate with another coat of black undercoat.

Unfortunately, the Q6 boiler steam test had to be abandoned due to poor steaming with only 30psi on the gauge, and time running out for the MPD boilersmith and fitter who finished their shift at 3pm. The two tea tray studs at that point were steam tight, with nothing leaking. The test was to resume the next day. The 9Fwas also in steam on Wednesday.

On Thursday (yesterday), Ian called in at the MPD about 12.30 to help out if necessary. However, there was only 15psi on the clock in spite of it being lit up at 06.30. By 14.00 it was showing 30psi and yet this was a warm boiler from the day before. So, again, the steam test was abandoned. It is thought that the tubes may be blocked so today, Friday, the MPD is going to blow through the tubes with compressed air before trying another steam test next Monday/Tuesday. Third time lucky?


2nd and 3rd March 2021

Two extremely cold and dull days at Grosmont this week for Bill Dobson, Nigel Bill, Gordon Wells, Jon Bradley and Ian Pearson on Tuesday, and Bill Dobson, Ed Bolam, Trevor Wilford and Ian Pearson on Wednesday.

63395 measuring up pipe runs for new lubrication system - Ian Pearson

Over the two days -

  1. Both pistons were removed from the J27.
  2. The route for the copper pipes for the new axlebox lubrication system were planned and measured (see photo).
  3. Copper pipe was cut to various lengths and labelled ready for soldering and bending (see photo).
  4. The eight new oil pipe brackets were reshaped.
  5. Five sections of backplate were rubbed down but it was too cold for them to be painted.
  6. Four taper pins were fitted to the brake hangers.
  7. A new steel pipe was fabricated for the ash pan spray system by Trevor Wilford.
  8. Both little end oil pots were removed to check the oil flow and refitted.
  9. Eight holes were drilled into the frames for fitting the new oil pipe brackets.

63395 copper pipe for new lubrication system cut and labelled on 3 March 2021 - Ian Pearson

At the MPD the 9F was having an insurance steam test.


23rd and 24th February 2021

Ian reports two rather mild windy days with an occasional shower at Grosmont this week, with Bill Dobson, Ian Pearson, Nigel Bill, Jon Bradley and Gordon Wells on Tuesday, and Bill Dobson, Ian Pearson and Ed Bolam on Wednesday.

The following jobs were done over the two days:

  1. Right hand connecting rod fitted to the crosshead and secured.
  2. Two new cotters fitted to the crosshead little end castellated nuts.
  3. Five sections of the cab backplate undercoated.
  4. Righthand piston fitted into the cylinder using the engine lift and lifting table.
  5. Lubricator, atomisers and lefthand piston rings loaded into Jon Bradley’s car for delivery to Matthew Storey's workshop for repairs and rectification, and new fittings on the atomisers.
  6. New taper pin split, cut, and fitted to the righthand piston head nut.
  7. Left and right little end oil pots fitted (see photo).
  8. New steel angle brackets cut and drilled to be fitted to the frames to enable the oil pipes to connect to the axle boxes (see photo).
  9. One of the brackets was temporarily fitted by Bill Dobson after drilling and tapping (see photo).
  10. Lefthand piston head nut cotter cut and fitted.
  11. Righthand piston packing fitted but not yet secured.

New angle brackets for fitting to frames to enable the oil pipes to connect to the axle boxes - Ian Pearson

One of the new angle brackkets for the oil pipes temporarily fitted after drilling and tapping on 24 February 2021 - Ian Pearson

Left hand oil pot fitted on 24 February 2021 - Ian Pearson

Elsewhere, as Chris Henwood has reported separately, there was a mega shunt at the MPD yesterday to get the J27 on to the back of No 4 Road so the buffer beam can be removed and work started on the cylinder rebore (see photo). A reminder therefore that there will be a J27 working party this Saturday 27 February to progress this work. If you are willing and able to help, please contact Chris direct on 07929 733934 - he would love to hear from you! On Saturday there will also be a working party on the K1 - please contact Paul Hutchinson direct on 07964 988551 if you want to join him.

65894 standing on No 4 Road at the MPD on 24 February 2021 to allow the buffer beam to be removed and the cylinders rebored - Ian Pearson

On the Railway itself, as many of you will be aware from Chris Price's recent video message, the NYMR is planning to reopen from 1 May.


16th and 17th February 2021

A fine, mild, day for a change at Grosmont on Tuesday for Bill Dobson, Nigel Bill, Jon Bradley and Ian Pearson, although the first job of the day remained lighting the stove.

All the lifting equipment was inspected, including 5 strops and one shackle from the K1 support coach. The only item to fail was the pull lift from Deviation Shed. This had numerous defects so has been scrapped. Nigel B is going to acquire a new one. The current colour code is now blue.

Bill and Gordon Wells spent all morning discussing and planning what type of fittings and components are required for the new Q6 lubrication system. Last Thursday, Ian went to Middlesbrough and picked up six 18 flexible pipes from Northern Hydraulics which will be used to feed the leading, intermediate, and driving axle boxes. Two 36” flexible pipes will be ordered for the trailing axle boxes shortly. It is intended to fit atomisers to the valves. Bill has acquired two old sets of atomisers but they will need testing for leaks. These were cleaned on the outside by Nigel B using the shot blaster at the MPD. It was also found that some of the threads which will connect the oil feeds have different threads. This will make fitting a bit more complicated and an adapter will have to be made. One of the plug fittings was missing - it was lost in the workshop some three years ago and thought to be somewhere underneath the workbench .

Jon primed the five sections of cab back plate fittings, and during the afternoon the left hand little end pin was fitted and secured.

Deviation Shed new clock from Ed Bolam - 17 February 2021 - Ian Pearson

Wednesday morning started a bit dull and overcast, albeit mild, but became much brighter later. Bill wasn't feeling very well so only Ian Pearson and Ed Bolam were present. The workshop clock had stopped working early last year (not a flat battery) and today Ed brought a new one and fitted it to the south end wall (see photo). Ian told Ed about the missing plug fitting from one of the atomisers which had been lost underneath the workbench. Ed bravely cleaned all the rubbish from under the work benches and eventually found the missing plug. He then continued to clean the stubborn dirt that was left on the atomisers'

body and threads. Ian spent most of the day putting a slight chamfer on the piston heads.

All the fire extinguishers had their yearly check - usually carried out in August but the company changed its routine check to February. Bill did call in with Zak the dog later in the afternoon as he was feeling better.


9th and 10th February 2021

Another misleading subject heading this week as Tuesday's working party had to be cancelled due to poor weather conditions. However, in spite of the poor road conditions, Ian Pearson and Bill Dobson made it to Grosmont on Wednesday, a bitterly cold day with the odd hail and snow showers. Not surprisingly, the first job was to get the stove lit and warmed up.

The left hand connecting rod fitted on 10 February 2021 with packing pending fitting of new bearing pads - Ian Pearson

The main job of the day was fitting both connecting rods using the new hydraulic lifting table. This made the job much easier than in the past, when lots of timber packing and manpower were required. Unfortunately the bearings could not be fitted, as new felt pads had had to be made for the little ends and big end bearings, and these are presently soaking in lubricating oil. Wooden packing pieces have therefore been put around the big end crank pins to keep the connecting rod from touching the crank pins (see photo). The little end pins and big end bearings will be fitted next week.

Bill has ordered 60 metres of copper pipe for the new axle box lubrication system using the 8 feed silvertown lubricator from the upstairs store in the workshop.


Page 3 of 34